OpenStudy (anonymous):

Solve this system. Please put your answer as an ordered pair. -x-7y=14 -4x-14y=-28

OpenStudy (abmon98):

* equation 1) by -4 \[-4(-x-7y)=14\]\[4x+28y=-56\] \[-4x-14y=-28\]\[28y+(-14y)=-56+(-28)\] \[14y=84\]\[y=6\] Plug that y value in any of the original 2 equations \[-x-7(6)=14\]\[-x=14+42\] \[x=-56\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (abmon98):

Your most welcome :)